OpenStudy (anonymous):

Mr.Atkinson wants to increase her circular garden by doubling the diameter. What effect will this have on the area?

OpenStudy (anonymous):

soo the answer 16

OpenStudy (anonymous):

can you show me how you got dat in equation

OpenStudy (ybarrap):

Area1= pi*r1^2, where r1 is radius of circle before increase of diameter Area2 = pi*(2*r1)^2, doubling diameter, doubles size of radius Area2=pi*2^2*r1^2, bring the factor "2" outside the exponent =4*pi*r1^2, since 2^2 is 4 =4*Area1, since Area1 = pi*r1^2 Thus, by doubling the diameter (r1), you double the radius (2*r1), which quadruples the initial Area (Area1).

OpenStudy (anonymous):

YBARRAP can u write that in equation i cant have writing soo can u show it in numbers and the answer please

OpenStudy (anonymous):

sorry - i made an error ybarrap is right - the radius iis doubled the area is increased 4 times pir^2 / pi (2r)^2 = r^2 / 4r^2 = 1 / 4

OpenStudy (anonymous):

so its 1/4

OpenStudy (anonymous):

well the ratio of smaller to larger is 1 / 4 thats the same as saying the larger one is 4 times bigger than the small (original) garden

OpenStudy (anonymous):

so what is the answerrrrr

OpenStudy (anonymous):

area increases 4 times

OpenStudy (anonymous):

okayy im confused it says what effect will this have on the area can u show me work and answer instead of wordsssss

OpenStudy (anonymous):

can u write it on the draw please

OpenStudy (anonymous):

ybarrap has shown you the work - and so have i

OpenStudy (anonymous):

ok - i'll try and draw it

OpenStudy (ybarrap):